# Dissipation-Driven Topological Phase Transition in Quantum Open Systems Independent of system Hamiltonian

### v1

@@ V Entanglement spectrum dynamics
Figure 4: Time-evolution of single particle entanglement spectrum. (a)(b) shows the single particle entanglement spectrum under periodical boundary conditions with gap closing dynamics and (c)(d) shows the single particle entanglement spectrum under open boundary conditions (OBC) with the topologically protected zero-mode structure. In (a)(c), the parameters is taken the same as Fig.2(a) and the time-evolving density matrix undergoes one topological phase transitions with its topological invariant changing from 1-1 to +1+1. In (b)(d), the parameters is taken the same as Fig.3(a) and the time-evolving density matrix undergoes two topological phase transitions with its topological invariant changing from $1$ to 1-1, and sequentially from 1-1 to 11. Also, we have taken $\hat{H_{0}}=0$ and the system size $N=500$ in this figure.
The characterization of topology through many-body entanglement spectrum analysis has been well-established in closed quantum systems [32, 33, 34]. Recent extensions to open quantum systems have revealed intriguing connections between entanglement spectrum crossings and non-Hermitian topology in Lindbladian dynamics[29]. Here, we establish a direct correspondence between the topology of mix states, encoded in their modular Hamiltonians, and the entanglement spectrum structure.
Entanglement spectrum is defined by the eigen values of $\hat{\rho}$. Given that $\hat{\rho}$ is a Gaussian state, the eigenvalue of $\hat{\rho}$, denoted by $\Xi_{\{\mathbf{e}\}}$ take the form of
$\Xi_{\{\mathbf{e}\}}=\prod_{i=1}^{N}\frac{1}{2}\left[1+(-1)^{e_{i}}\epsilon_{i}\right].$ (35)
Here, $e_{i}$ takes values 0 or $1$ and the notation $\{\mathbf{e}\}$} denotes the complete set of all possible configurations of $e_{i}$. And $\epsilon_{i}$ is the eigen values of $\Delta$, usually named as the single-particle entanglement specreum (SPES). The single-particle entanglement spectrum not only fully determines the many-body entanglement spectrum, but also serves as a direct indicator of the topological characteristics inherent in time-evolved density operators $\hat{\rho}(t)$.
@@ VI Summary
We firstly employ a vectorization procedure for Eq. (14), which is transformed into
@@
$|\Delta_{s}\rangle=i4[(\sum_{m=\pm}\lambda_{m}|\psi_{m}\rangle\langle\chi_{m}|)\otimes(\sum_{n=\pm}|\psi_{n}^{*}\rangle\langle\chi_{n}^{*}|)$
$-(\sum_{m=\pm}|\psi_{m}\rangle\langle\chi_{m}|)\otimes(\sum_{n=\pm}\lambda_{n}^{*}|\psi_{n}^{*}\rangle\langle\chi_{n}^{*}|)]^{-1}|Y\rangle$
$=\sum_{m,n=\pm}\frac{1}{(\lambda_{m}-\lambda_{n}^{*})}|\psi_{m}\rangle\langle\chi_{m}|\otimes|\psi_{n}^{*}\rangle\langle\chi_{n}^{*}|]|Y\rangle,$ (38)
where we have used the completeness relation $\sum_{m=\pm}|\psi_{m}\rangle\langle\chi_{m}|$. Rewriting Eq. (A) in the form of matrix, Eq. (15) is obtained.
@@
Acknowledgements This work was supported by the National Natural Science Foundation of China (grant NO. 12204406), the Yunnan Fundamental Research Projects (grant NO. 202401CF070187), and the Start-up Research Fund of Southeast University (RF1028624190)..

### v2

@@ V Entanglement spectrum dynamics
Figure 4: Time-evolution of single particle entanglement spectrum. (a)(b) shows the single particle entanglement spectrum under periodical boundary conditions with gap closing dynamics and (c)(d) shows the single particle entanglement spectrum under open boundary conditions (OBC) with the topologically protected zero-mode structure. In (a)(c), the parameters is taken the same as Fig.2(a) and the time-evolving density matrix undergoes one topological phase transitions with its topological invariant changing from 1-1 to +1+1. In (b)(d), the parameters is taken the same as Fig.3(a) and the time-evolving density matrix undergoes two topological phase transitions with its topological invariant changing from $1$ to 1-1, and sequentially from 1-1 to 11. Also, we have taken the system Hamiltonian of Eq.(1) as $\hat{H}=J\hat{c}_{m}^{\dagger}\hat{c}_{m+1}+J\hat{c}_{m+1}^{\dagger}\hat{c}_{m}$ with $J/v_{1}=1$ and the system size $N=500$ in this figure.
The characterization of topology through many-body entanglement spectrum analysis has been well-established in closed quantum systems [32, 17, 39]. Recent extensions to open quantum systems have revealed intriguing connections between entanglement spectrum crossings and non-Hermitian topology in Lindbladian dynamics[38]. Here, we establish a direct correspondence between the topology of mixed states, encoded in their modular Hamiltonians, and the entanglement spectrum structure.
For a Gaussian state, the entanglement spectrum, defined by the eigen values of density matrix $\hat{\rho}$, is fully determined by the correlation matrix $\Delta(t)$. The computation proceeds by first diagonalizing $\Delta(t)$ to obtain its eigenvalues, known as the single-particle entanglement spectrum (SPES) $\epsilon_{i}$}, which results in a set of real eigenvalues due to the matrix’s Hermiticity. Subsequently, the full many-body entanglement spectrum $\Xi_{{\mathbf{e}}}$ is built from these single-particle eigenvalues via the relation
$\Xi_{\{\mathbf{e}\}}=\prod_{i=1}^{N}\frac{1}{2}\left[1+(-1)^{e_{i}}\epsilon_{i}\right],$ (35)
where ${\mathbf{e}}=(e_{1},e_{2},\dots,e_{N})$ with $e_{i}\in\{0,1\}$ denotes the occupation configuration of the single-particle modes[17]. In our numerical simulations, we compute $\Delta(t)$ at each time step by solving Eq.(9) and subsequently diagonalize it to obtain the SPES. This procedure ensures that the topological features of the state are directly accessible from the SPES.
@@ VI Summary
We firstly employ a vectorization procedure[14]for the steady-state equation for the correlation matrix Eq. (14). The vectorization map $\text{vec}(\cdot)$ transforms a matrix into a column vector by stacking its columns. A fundamental property of this map is:
$\text{vec}(A\rho B)=(A\otimes B^{T})\text{vec}(\rho),$ (36)
which, in the Dirac notation used in our manuscript ($|\rho\rangle\equiv\text{vec}(\rho)$), becomes:
$A\rho B\rightarrow(A\otimes B^{T})|\rho\rangle.$ (37)
Applying this property term by term to our steady-state equation:
$H_{\text{eff}}\Delta_{s}\rightarrow(H_{\text{eff}}\otimes I)|\Delta_{s}\rangle,$
$\Delta_{s}H_{\text{eff}}^{\dagger}\rightarrow(I\otimes(H_{\text{eff}}^{\dagger})^{T})|\Delta_{s}\rangle=(I\otimes H_{\text{eff}}^{*})|\Delta_{s}\rangle,$
$4iY\rightarrow 4i|Y\rangle,$ (38)
then Eq. (14) is transformed into
@@
$|\Delta_{s}\rangle=i4[(\sum_{\alpha=\pm}\lambda_{\alpha}|\psi_{\alpha}\rangle\langle\chi_{\alpha}|)\otimes(\sum_{\beta=\pm}|\psi_{\beta}^{*}\rangle\langle\chi_{\beta}^{*}|)$
$-(\sum_{\alpha=\pm}|\psi_{\alpha}\rangle\langle\chi_{\alpha}|)\otimes(\sum_{n=\pm}\lambda_{\beta}^{*}|\psi_{\beta}^{*}\rangle\langle\chi_{\beta}^{*}|)]^{-1}|Y\rangle$
$=\sum_{\alpha,\beta=\pm}\frac{1}{(\lambda_{\alpha}-\lambda_{\beta}^{*})}|\psi_{\alpha}\rangle\langle\chi_{\alpha}|\otimes|\psi_{\beta}^{*}\rangle\langle\chi_{\beta}^{*}|]|Y\rangle,$ (41)
where we have used the completeness relation $\sum_{\alpha=\pm}|\psi_{\alpha}\rangle\langle\chi_{\alpha}|$. Rewriting Eq. (A) in the form of matrix, Eq. (15) is obtained.
@@
Acknowledgements This work was supported by the National Natural Science Foundation of China (grant NO. 12204406 and Grand No. 12504187), the Yunnan Fundamental Research Projects (grant NO. 202401CF070187), and the Start-up Research Fund of Southeast University (RF1028624190).
Code availability. Codes for numerical calculations and figures are available at [44].